we have 3 blue and 5 red balls lay in a box three balls are chosen randomly out of the box find the probability distribution for number of blue balls without replacement.


1
Expert's answer
2021-05-25T13:11:02-0400

Let B denotes Blue ball

and R denotes Red ball


Let X denotes the event of number of blue balls are chosen out of three balls drawn from the box


So,

Case 1:

When X = 0 means all three balls drawn are Red

"P(X_0)=\dfrac{5}{8}\times \dfrac{4}{7}\times\dfrac{3}{6}=\dfrac{60}{336}"



Case2:

When X=1 means only one Blue balls are chosen out of three balls

and it can be BRR + RBR +RRB

"So, P(X_1)=(\dfrac{3}{8}\times\dfrac{5}{7}\times\dfrac{4}{6})+(\dfrac{5}{8}\times\dfrac{3}{7}\times\dfrac{4}{6})+(\dfrac{5}{8}\times\dfrac{4}{7}\times\dfrac{3}{6})=\dfrac{180}{336}"



Case3:

When X=2 means two blue balls are there on choosing 3 balls from the box

and it can be BBR +BRB+ RBB

So, "P(X_2)=(\dfrac{3}{8}\times\dfrac{2}{7}\times\dfrac{5}{6})+(\dfrac{3}{8}\times\dfrac{5}{7}\times\dfrac{2}{6})+(\dfrac{5}{8}\times\dfrac{3}{7}\times\dfrac{2}{6})=\dfrac{90}{336}"



CAse4:

When X=3 means all blue balls are chosen from the box

So, "P(X_3)=\dfrac{3}{8}\times\dfrac{2}{7}\times\dfrac{1}{6}=\dfrac{6}{336}"


So, Probability distribution table will be: